OpenStudy (anonymous):

flying into the wind, a plane take 6h to fly 3000 km. On the return flight, with the same wind, the plane takes 5 h to complete the trip. How fast does the plane fly without any wind, and how fast was the wind blowing?

OpenStudy (anonymous):

3000 = (p-w)6 3000 = (p+w)5 p is plane speed w is wind speed two equations and two unknowns. solve for w

OpenStudy (anonymous):

3000/6 = p-w 3000/5 = p+w 2w = 100 w = 50 kmph

OpenStudy (anonymous):

let speed of wind=w and speed of plane alone=p now we know ,distance=speed x time let distance=x (w+p) x 6 hours=3000 km and (p-w) x 5 hours =3000 km 6w+6p=3000=5p-5w p=3000=-11w lets take p=-11w 5(p-w)=3000 5(-12w)=3000 -60w=3000 w=50 km/hr